# W-HMR: Human Mesh Recovery in World Space with Weak-supervised Camera Calibration and Orientation Correction

### 29 Nov 2023 v1

@@ Abstract
For a long time, in the field of reconstructing 3D human bodies from monocular images, most methods opted to simplify the task by minimizing the influence of the camera. Using a coarse focal length setting results in the reconstructed bodies not aligning well with distorted images. Ignoring camera rotation leads to an unrealistic reconstructed body pose in world space. Consequently, existing methods’ application scenarios are confined to controlled environments. And they struggle to achieve accurate and reasonable reconstruction in world space when confronted with complex and diverse in-the-wild images. To address the above issues, we propose W-HMR, which decouples global body recovery into camera calibration, local body recovery and global body orientation correction. We design the first weak-supervised camera calibration method for body distortion, eliminating dependence on focal length labels and achieving finer mesh-image alignment. We propose a novel orientation correction module to allow the reconstructed human body to remain normal in world space. Decoupling body orientation and body pose enables our model to consider the accuracy in camera coordinate and the reasonableness in world coordinate simultaneously, expanding the range of applications. As a result, W-HMR achieves high-quality reconstruction in dual coordinate systems, particularly in challenging scenes. Codes will be released on project page after publication.
1 Introduction
Recovering 3D human meshes from monocular images is a crucial task in human-centered research, with wide-ranging applications in areas such as try-on, animation, and virtual reality. However, there is a noticeable lack of research on global body reconstruction from images captured under complex shooting conditions. The reconstructed bodies outside controlled environments are often unsuitable for real-world applications, as depicted in Fig. 2 and Fig. 5.
@@ 1 Introduction
The primary challenge in this task arises from the limited camera information. Among camera parameters, focal length plays a critical role in distortion effects, influencing the alignment of reconstructed 3D human bodies with original images. Additionally, camera rotation determines the reasonableness of the reconstructed body pose in the world (global) or camera (local) coordinate after coordinate transformation. Fig. 1 illustrates the treatment of the camera in existing models, emphasizing the need for improved methods. The focal length either involves simple manual definitions as seen in HMR [14] and CLIFF [25], or network predictions as utilized by SPEC [19] and Zolly [43]. As for camera rotation, most models [52, 53, 51, 20, 17, 18, 26] do not consider it and recover body in camera coordinate, except for SPEC. Even based on the most comprehensive camera model, SPEC still has limitations due to the inherent flaws in its CamCalib model.
Firstly, the focal length is predicted based on background environmental information, which is inadequate for the body recovery task and cannot handle finer distortions of body parts. To address these challenges, we propose a weak-supervised camera calibration method. The core idea is that we don’t need to predict the original focal length. This method leverages body distortion information to predict a reasonable focal length, eliminating the need for supervision with rare focal length labels and enhancing perspective projection accuracy. Based on the full-perspective camera model, 2D supervision becomes more precise, thereby improving recovery accuracy.
Secondly, the predicted camera rotation is not entirely accurate. With pseudo camera rotation as input, training the model like SPEC to predict poses in world coordinate would lead to error accumulation. The error in pseudo camera parameters would impact the final recovery precision. Consequently, the accuracy of recovery is limited. Additionally, it decreases the model’s generalizability. For example, SPEC’s performance would be worse when the camera rotation is known, as shown in Tab 1. We adopt an indirect approach for predicting body poses in world coordinate. First, we train the model to achieve optimal accuracy of pose and shape in camera coordinate. Then, we propose a novel module named OrientCorrect to assist in correcting body orientation, thus avoiding the issues above.
@@
    We present the first weak-supervised camera calibration method focus on the human recovery task. We achieve reasonable focal length prediction without focal length labels, but relying on human body distortion information. Better mesh-image alignment makes 2D supervision more precise and effectively improves recovery accuracy.
@@
    We design a practical orientation correction module. The module solves the issue that previous methods cannot generate reasonable human poses in world space. This method is simple and effective, and can be well transplanted to other models without destroying the original structure and performance.
@@
    Our decoupling idea successfully mitigates the negative impact of incorrect camera parameter estimation. As a result, our model can achieve remarkable reconstruction results in both coordinate systems simultaneously, which is not possible with past models.
@@ 2.1 Camera Model in Human Recovery
To the best of our knowledge, Kissos et al. [16] was the first to enhance the weak-perspective camera model in pose estimation. They assumed a constant camera Field of View (FoV) of 55 degrees instead of a fixed focal length. Later, CLIFF [25] also adopted the same camera model. SPEC [19] took a different approach by constructing a new dataset and training a camera parameter prediction model named CamCalib. Zolly [43] predicted the focal length indirectly through a carefully designed translation prediction model, albeit with partial reliance on focal length labels and limitations in focal length prediction for cropped images. Regarding camera rotation, it was usually overlooked in previous works. SPEC was the pioneering work to advocate for the recovery of the body in world coordinate for postural plausibility. It relied on CamCalib to predict camera parameters and used these parameters as model inputs to directly predict the global body. However, its one-step approach led to error accumulation, constraining reconstruction accuracy and application scope. In contrast, our decoupling of this process effectively addresses these issues.
@@ 2.2 Regression-based Method
In the domain of 3D human body reconstruction, methods can be broadly classified into regression-based and optimization-based approaches. Optimization-based methods are beyond the scope of this paper and can be explored in the literature review [40]. Regression-based models can be further categorized into model-based and model-free approaches based on the output type. Model-based models refer to those that output template parameters, with the majority of current models [14, 8, 52, 17, 18] still relying on the SMPL algorithm [29], and a few [37, 57, 49, 50] based on SCAPE [2], GHUM&GHUML [46], and SMPL-X [36]. These models leveraged the prior knowledge embedded in parameter models but were limited to reconstructing finer mesh details due to constraints in degrees of freedom.
In contrast, model-free models in this field output voxels [41, 56], UV position maps [54, 39] and 3D vertex coordinates [26, 33, 27, 5]. Here, we mainly focus on methods that output 3D vertex coordinates. Apart from methods that output voxels, all other output types would be converted to 3D vertices and then transformed into meshes based on predefined triangle faces. One classic model-free method, Graphormer [26], designed a regression module based on transformer[6], which directly outputs 431 3D sparse vertices. The advantage of model-free models is that they can recover more details than model-based models. However, their limitations lie in their inability to effectively utilize prior knowledge of the human body, and their models tend to have larger sizes. To overcome the limitations of both model-based and model-free approaches, we propose a hybrid method where the model-based regression module provides a strong initialization, and the model-free module further refines the results. This approach aims to achieve optimal mesh precision with relatively fewer parameters, bridging the gap between model-based and model-free methods.
@@ 3.1 Preliminary
The whole process is how a 3D object is imaged on a 2D plane based on the perspective camera model. You may notice from Equ. 2 that when $f$ is set very large, the effect of $Z$ would become negligible and the perspective effect disappears. The consequence is that the human mesh does not align well with the original image, making 2D supervision in error and ultimately affecting recovery accuracy.
Translation and Scale Following  [14], we continue to utilize SMPL parameters $\varTheta$ as our template. The SMPL parameters encompass pose $\boldsymbol{\theta}\in\mathbb{R}^{24\times 3}$, shape $\boldsymbol{\beta}\in\mathbb{R}^{10}$, where $\boldsymbol{\theta}_{0}$ is the body orientation. Specifically, pose $\boldsymbol{\theta}$ and shape $\boldsymbol{\beta}$ can be inputted into a pre-defined function $M\left(\boldsymbol{\theta},\boldsymbol{\beta}\right)\in\mathbb{R}^{3\times N}$ to obtain $N$ vertex coordinates, i.e., $(X,Y,Z)$ we talk about before. The mesh derived from these vertices is the reconstructed human mesh. To project it onto the image, the model outputs three additional parameters, i.e., translation $(t_{x},t_{y})$, and scale $\boldsymbol{s}$. As for $\boldsymbol{s}$ and $(t_{x},t_{y})$, they serve to determine the size and position of the human meshes within the cropped image. Based on the usual assumptions, we assume a human body range of 2 meters, which establishes the relationship between $t_{b}^{z}$ and $f$ as
@@ 3.2 Weak-supervised Camera Calibration
Focal length prediction is essentially an ill-posed problem. From Sec. 3.1, we can see that there are too many influencing factors in the process from 3D objects to 2D images. Trying to predict the true focal length through a neural network is almost impossible. So, we propose to use the human body distortion information to predict a ”reasonable” focal length for full-perspective projection.
Camera Calibration The focal length predictor takes the final feature map $\varPhi_{3}$ as input and output a single value $t_{b}^{z}$. To constrain the range of the predicted $t_{b}^{z}$ [43], we employ a sigmoid function and scale it by 10. This adjustment is necessary because distortion becomes irrelevant when a person stands 10 meters from the camera. We add batch normalization to avoid gradient vanishing or exploding caused by the sigmoid function. The focal length can be calculated according to $f=\frac{2sh_{bbox}t_{b}^{z}}{2}$, where the height of bounding box $h_{bbox}$ is known and scale $s$ is output by the PyMAF-like regression module. From Equ. 2, to implement the projection, we also need to know the translation of the body to the camera. $t_{b}^{z}$ has already been generated, so we need to further compute $(t_{b}^{x},t_{b}^{y})$. Following [16], $(t_{b}^{x},t_{b}^{y})=(t_{x},t_{y})$ when projecting in cropped image. But in the full image, the conversion functions for $(t_{x},t_{y})$ to $(t_{x}^{b},t_{y}^{b})$ are summarized below
@@
where $(c^{x}_{bbox},\,c^{y}_{bbox})$ is the center of the bounding box and $(W,H)$ is the width and height of the full image. The reason why projection is important in the human recovery task is that we need to use rich 2D joint labels to enhance supervision and improve model performance.
Weak Supervision Our main losses can be categorized into three kinds: joint loss, vertex loss, and parameter loss. In this section, we only introduce the joint loss. Please refer to the supplementary material for information on the others. The joint loss can be calculated using the following function:
@@
Figure 3: This figure depicts our model training paradigm, illustrating the detachment strategy at different stages to achieve stable model training. The colorful symbol $\otimes$ represents detachment. Traditional 2D supervision refers to projecting 3D joints onto the cropped image to get 2D joints based on the weak-perspective camera model, just as traditional models would do.
@@ 3.3 Orientation Correction
Our goal is to recover human mesh in world space. Weak-supervised camera calibration can only help the model to get the optimal mesh accuracy in camera coordinate. After that, we need to consider how to keep the human pose reasonable in world coordinate without destroying the original accuracy. In short, we need to predict the body orientation in world coordinate to replace the body orientation generated in camera coordinate.
@@
$R_{b}=R_{b}^{c}+FC\left(\oplus\left(\phi_{i}^{s},\,\,glob\_info,R_{c},R_{b}^{c}\right)\right)$ (6)
@@
where $\oplus$ represents concatenation, $FC(\cdot)$ are couples of FC layers. The global information $glob\_info=\left\{\frac{\varDelta c_{x}}{\tilde{f}},\frac{\varDelta c_{y}}{\tilde{f}},\frac{h_{bbox}}{\tilde{f}},\frac{W}{\tilde{f}},\frac{H}{\tilde{f}}\,\,\right\}$ and it is important auxiliary information for predicting body orientation [25]. As for $\tilde{f}$, it is for normalization and $\tilde{f}=\sqrt{W^{2}+H^{2}}$. This iterative approach is employed to obtain the final body orientation in world coordinate. The OrientCorrect is a completely separate module. But it not only makes full use of what the model has learnt in camera coordinate, but also avoids destroying the body feature space that the model has constructed. Additionally, this method can be easily applied to other models based on camera coordinate without reducing the original precision.

### 25 Mar 2024 v3

@@ Abstract
For a long time, in reconstructing 3D human bodies from monocular images, most methods opted to simplify the task by minimizing the influence of the camera. Using a coarse focal length setting results in the reconstructed bodies not aligning well with distorted images. Ignoring camera rotation leads to an unrealistic reconstructed body pose in world space. Consequently, the application scenarios of existing methods are confined to controlled environments. When confronted with complex and diverse in-the-wild images, they struggle to achieve accurate and reasonable reconstruction in world space. To address the above issues, we propose W-HMR, which decouples global body recovery into camera calibration, local body recovery, and global body orientation correction. We design the first weak-supervised camera calibration method for body distortion, eliminating dependence on focal length labels and achieving finer mesh-image alignment. We propose a novel orientation correction module to allow the reconstructed human body to remain normal in world space. Decoupling body orientation and body pose enables our model to consider the accuracy in camera coordinate and the reasonableness in world coordinate simultaneously, expanding the range of applications. As a result, W-HMR achieves high-quality reconstruction in dual coordinate systems, particularly in challenging scenes. Codes and demos have been released on the project page https://yw0208.github.io/w-hmr/.
keywords:
Human mesh recovery, monocular 3D reconstruction, world space, camera calibration, motion capture
[…]
Qualitative results on SPEC-MTP [19]. The images in each triplet from left to right are the input image, the output in the camera coordinate, and the output in world coordinate. Compared to traditional methods, W-HMR shows obvious advantages for human mesh recovery in world space. W-HMR can efficiently correct erroneous poses in camera coordinate and ensure the reasonableness of poses in world space.
Introduction
Recovering 3D human meshes from monocular images is crucial in human-centered research, with wide-ranging applications in areas such as try-on, animation, and virtual reality. However, there is a noticeable lack of research on global body reconstruction from images captured under complex shooting conditions. For previous methods based on camera coordinate, the reconstructed bodies outside controlled environments are often unsuitable for real-world applications, as depicted in W-HMR: Human Mesh Recovery in World Space with Weak-supervised Camera Calibration and Orientation Correction.
@@ Introduction
The primary challenge in this task arises from the limited camera information. Among camera parameters, focal length is critical in distortion effects, influencing the alignment of reconstructed 3D human bodies with original images. There is nothing inherently wrong with this misalignment. However, since most existing models are trained with 2D joint labels, this misalignment due to distortion affects the training and ultimately leads to decreased recovery accuracy. Additionally, camera rotation determines the reasonableness of the reconstructed body pose in the world (global) or camera (local) coordinate after coordinate transformation. Fig. 1 illustrates the treatment of the camera in existing models, emphasizing the need for improved methods. The focal length either involves simple manual definitions as seen in HMR [14] and CLIFF [25], or network predictions as utilized by SPEC [19] and Zolly [43]. As for camera rotation, most models [54, 53, 52, 20, 17, 18, 27] do not consider it and recover body in camera coordinate, except for SPEC. Even based on the most comprehensive camera model, SPEC still has limitations due to the inherent flaws in its CamCalib model, which is discussed as follows.
Firstly, according to traditional methods like SPEC, the focal length is predicted based on background environmental information, which cannot handle finer distortions of body parts and is inadequate for the body recovery task. Moreover, focal length labels are always rare and not diverse enough, which can’t fit the need for model training. To address these challenges, we propose a weak-supervised camera calibration method. The core idea is that we don’t need to predict original focal lengths but the "reasonable" ones. Our method leverages body distortion information to predict a reasonable focal length for perspective projection, eliminating the need for supervision with rare focal length labels and making it more suitable for body tasks. Based on the full-perspective camera model, 2D supervision becomes more precise, thereby improving recovery accuracy.
Secondly, the predicted camera rotation is not entirely accurate. With pseudo camera rotation as input, training a model like SPEC to predict poses in world coordinate would lead to error accumulation. The error in pseudo camera parameters would accumulate during training and impact the final recovery precision. Consequently, the accuracy of recovery is limited. Additionally, it decreases the model’s generalizability. For example, compared to SPEC’s advantages in world space, it performs poorly in camera coordinate, as shown in Table 2. We adopt an indirect approach for predicting body poses in world coordinate. First, we train our model to achieve optimal pose and shape accuracy in camera coordinate. Then, we propose a novel module named OrientCorrect to assist in correcting body orientation, thus avoiding the issues above and enabling human mesh recovery in world coordinate.
@@
    We present the first weak-supervised camera calibration method focus on the human recovery task. We achieve reasonable focal length prediction without focal length labels, but relying on human body distortion information. Better mesh-image alignment makes 2D supervision more precise and effectively improves recovery accuracy. Our method even outperforms the latest supervised method like Zolly [43].
@@
    We design a practical orientation correction module. The module solves the issue that previous methods cannot generate reasonable human poses in world space. This method is simple and effective and can be well transplanted to other models without destroying the original structure and performance.
@@
    Our decoupling idea successfully mitigates the negative impact of incorrect camera parameter estimation. As a result, our model can achieve remarkable reconstruction results in both coordinate systems simultaneously, which is impossible with past models.
@@ 1.1 Camera Model in Human Recovery
To the best of our knowledge, Kissos et al. [16] was the first to enhance the weak-perspective camera model in pose estimation. They assumed a constant camera Field of View (FoV) of 55 degrees instead of a fixed focal length. Later, CLIFF [25] also adopted the same camera model. SPEC [19] took a different approach by constructing a new dataset and training a camera parameter prediction model named CamCalib. The latest method, Zolly [43], indirectly predicts the focal length through a well-designed translation prediction module, albeit partially relying on focal length labels and only predicting the focal length of cropped images. Regarding camera rotation, it was usually overlooked in previous works. SPEC was the pioneering work to advocate for the recovery of the body in world coordinate for postural plausibility. It relied on CamCalib to predict camera parameters and used these parameters as model inputs to directly predict the global body. However, its one-step approach led to error accumulation, constraining reconstruction accuracy and application scope. In contrast, our decoupling of this process effectively addresses these issues.
@@ 1.2 Regression-based Method
In the 3D human body reconstruction domain, methods can be broadly classified into regression-based and optimization-based approaches. Optimization-based methods are beyond the scope of this paper and can be explored in the literature review [40]. Regression-based models can be further categorized into model-based and model-free approaches based on the output type. Model-based models refer to those that output template parameters, with the majority of current models [14, 8, 54, 17, 18] still relying on the SMPL algorithm [29], and a few [37, 58, 50, 51] based on SCAPE [2], GHUM&GHUML [47], and SMPL-X [36]. These models leveraged the prior knowledge embedded in parameter models but were limited to reconstructing finer mesh details due to constraints in degrees of freedom.
In contrast, model-free models in this field output voxels [41, 57], UV position maps [55, 39] and 3D vertex coordinates [27, 34, 26, 5]. Here, we mainly focus on methods that output 3D vertex coordinates. Apart from methods that output voxels, all other types of output would be converted to 3D vertices and then transformed into meshes based on predefined triangle faces. One classic model-free method, Graphormer [27], designed a regression module based on transformer [6], which directly outputs 431 3D sparse vertices. The advantage of model-free models is that they can recover more details than model-based models. However, their limitations lie in their inability to effectively utilize prior knowledge of the human body, and their models tend to have larger sizes. To overcome the limitations of both model-based and model-free approaches, we propose a hybrid method where the model-based regression module provides a strong initialization, and the model-free module further refines the results. This approach aims to achieve optimal mesh precision with relatively fewer parameters, bridging the gap between model-based and model-free methods.
[…]
@@ 2 Method
The overview of W-HMR is shown in Fig. 2. The input image undergoes a PyMAF-like model to extract features and obtain the recovered human mesh in camera coordinates. We design a focal length predictor to predict the focal length implementing a full-perspective projection on the full image for 2D joint supervision. In addition, we propose the OrientCorrect module to correct the body orientation to keep the pose reasonable in the world space. Our method is introduced in five sections. We first present some prior knowledge in Section 2.1. This explains why we need to use the full perspective camera model and the full image information. In Section 2.2, we focus on how we predict the focal length and how we use the predicted focal length for supervision. In Section 2.3, we describe how OrientCorrect works, which is another important module. In Section 2.4, we describe how to stabilize model training without focal length labels. Finally, in Section 2.5, we add some introduction to more loss functions. To fully demonstrate the effectiveness of our method, we experiment with various network structures. Specific details can be found in Section 3.1.
@@ 2.1 Preliminary
The whole process is how a 3D object is imaged on a 2D plane based on the perspective camera model. You may notice from Equ. 2 that when $f$ is set very large, the effect of $Z$ would become negligible and the perspective effect would disappear. The consequence is that the human mesh does not align well with the original image, making 2D supervision in error and ultimately affecting recovery accuracy.
Translation and Scale Following  [14], we continue to utilize SMPL parameters $\varTheta$ as our template. The SMPL parameters encompass pose $\boldsymbol{\theta}\in\mathbb{R}^{24\times 3}$, shape $\boldsymbol{\beta}\in\mathbb{R}^{10}$, where $\boldsymbol{\theta}_{0}$ is the body orientation. Specifically, pose $\boldsymbol{\theta}$ and shape $\boldsymbol{\beta}$ can be inputted into a pre-defined function $M\left(\boldsymbol{\theta},\boldsymbol{\beta}\right)\in\mathbb{R}^{N\times 3}$ to obtain $N$ vertex coordinates, i.e., $(X,Y,Z)$ we talked about before. The mesh derived from these vertices is the reconstructed human mesh. To project it onto the image, the model outputs three additional parameters, i.e., translation $(t_{x},t_{y})$, and scale $\boldsymbol{s}$. As for $\boldsymbol{s}$ and $(t_{x},t_{y})$, they determine the size and position of the human meshes within the cropped image. Based on the usual assumptions, we assume a human body range of 2 meters, which establishes the relationship between $t_{b}^{z}$ and $f$ as
@@ 2.2 Weak-supervised Camera Calibration
Focal length prediction is essentially an ill-posed problem. From Section 2.1, we can see too many influencing factors in the process from 3D objects to 2D images. Trying to predict the actual focal length through a neural network is almost impossible. So, we propose using human body distortion information to predict a "reasonable" focal length for full-perspective projection.
Camera Calibration The focal length predictor takes the final feature map $\varPhi_{3}$ as input and output a single value $t_{b}^{z}$. To constrain the range of the predicted $t_{b}^{z}$ [43], we employ a sigmoid function and scale it by 10. This adjustment is necessary because distortion becomes irrelevant when a person stands 10 meters from the camera. We add batch normalization to avoid gradient vanishing or exploding caused by the sigmoid function. The focal length can be calculated according to $f=\frac{sh_{bbox}t_{b}^{z}}{2}$, where the height of bounding box $h_{bbox}$ is known and scale $s$ is output by the PyMAF-like regression module. From Equ. 2, to implement the projection, we also need to know the translation of the body to the camera. $t_{b}^{z}$ has already been generated, so we need further to compute $(t_{b}^{x},t_{b}^{y})$. Following [16], $(t_{b}^{x},t_{b}^{y})=(t_{x},t_{y})$ when projecting in cropped image. But in the full image, the conversion functions for $(t_{x},t_{y})$ to $(t_{x}^{b},t_{y}^{b})$ are summarized below
@@
where $(c^{x}_{bbox},\,c^{y}_{bbox})$ is the center of the bounding box and $(W,H)$ is the width and height of the full image. The reason why projection is important in the human recovery task is that we need to use rich 2D joint labels to enhance supervision and improve model performance. By the way, we also explain why we keep emphasizing the full image here. For example, as shown in Fig. 3, the distortion of stretching to left and right in red circles can only be achieved by projecting on the full image rather than solely on cropped images. That’s why we emphasize the implementation of full perspective projection in the full image. Moreover, the actual body orientation cannot be correctly reflected in the cropped image when the human is not in the center of the original image. Therefore, full-image information, i.e., global information in our method, is also an important part of our model input.
Figure 3: That image on the left is the original image. The middle shows the cropped images used in the traditional method at the top. The predicted focal length, body orientation, and shooting angle in traditional methods are shown at the bottom of the middle image. The image on the right shows the actual shot, illustrating that cropping leads to the loss of necessary information. This loss of necessary information ultimately affects the reconstruction results.
Weak Supervision Our main losses can be categorized into joint loss, vertex loss, and parameter loss. In this section, we only introduce the joint loss. Please refer to Section 2.5 for information on the others. The joint loss can be calculated using the following function:
@@
Figure 4: This figure depicts our model training paradigm, illustrating the detachment strategy at different stages to achieve stable model training. The symbol $\otimes$ represents detachment. Traditional 2D supervision refers to projecting 3D joints onto the cropped image to get 2D joints based on the weak-perspective camera model, just as traditional models did. The "freeze" means other modules are freezed in the third stage and only OrientCorrect are updated.
@@ 2.3 Orientation Correction
Our goal is to recover human meshes in world space. Weak-supervised camera calibration can only help the model to get the optimal mesh accuracy in camera coordinate. After that, we need to consider how to keep the human pose reasonable in world coordinate without destroying the original accuracy. In short, we need to predict the body orientation in world coordinate to replace the body orientation predicted in camera coordinate.
@@
$R_{b}=R_{b}^{c}+FC\left(\oplus\left(\phi_{3}^{s},\,glob\_info,\,R_{c},\,R_{b}^{c}\right)\right)$ (6)
@@
where $\oplus$ represents concatenation, $FC(\cdot)$ are couples of FC layers. To avoid distress, we explain why the operation of rotation matrices here does not use multiplication but rather direct addition. The network of OrientCorrect is similar to the classic iterative regressor used in HMR [14]. The similar design, instead of multiplication, has been employed in many previous works [25, 18, 54, 44] to predict rotations. The global information $glob\_info=\left\{\frac{\varDelta c_{x}}{\tilde{f}},\frac{\varDelta c_{y}}{\tilde{f}},\frac{h_{bbox}}{\tilde{f}},\frac{W}{\tilde{f}},\frac{H}{\tilde{f}}\,\,\right\}$ and it is important auxiliary information for predicting body orientation [25]. As for $\tilde{f}=\sqrt{W^{2}+H^{2}}$ in global information, it is just for normalization here and does not hold any geometric meanings, so we don’t use the predicted focal length here. The OrientCorrect is a completely separate module. So, it not only makes full use of what the model has learned in camera coordinate, but also avoids destroying the body feature space that the model has constructed. Additionally, this method can be easily applied to other models based on camera coordinate without reducing the original precision, which is shown in Table 8.
2.4 Training Paradigm
[…]
$\mathcal{L}_{f}=MSE\left(f,\tilde{f}\right)$ (7)
where $\tilde{f}=\sqrt{W^{2}+H^{2}}$ serves as pseudo focal length labels to ensure focal length predictor outputs coarse but relatively reasonable results. Note it is only applied in stage I.
[…]
In the first two stages, training is conducted entirely in camera coordinate. In the third stage, we need to equip the model to reconstruct the human body in world coordinate. The pre-trained regressors, focal length predictor, and other modules are frozen at this stage. We only need to train the OrientCorrect, which is composed of three FC layers. We exclusively use the 3D joint, mesh vertices, and SMPL parameters in world coordinate for supervision.
2.5 Other Losses
When we employ a hybrid regression module, we obtain 431 sparse vertices, 1723 coarse vertices, and 6840 complete vertices from the vertex regressor. Therefore, we design a triple vertex loss function, denoted as
where $\beta$ is shape parameters in SMPL and $\theta$ is pose parameters represented as rotation matrices.
To enhance the model’s ability to focus on human regions and extract refined body information, we include two additional sub-tasks for auxiliary supervision, which are inspired by [54].
First, let’s delve into the IUV map, which is integral to the IUV loss. The IUV map, as defined in [10], uses "I" to indicate the body part to which a pixel belongs and "UV" to represent the 2D coordinates on the surface of the body part. This mapping establishes dense correspondences from 3D mesh to 2D plane. We employ three 2D convolution layers to predict the IUV map based on the feature map $\varPhi_{3}$ and render the ground truth online. The loss function is defined as
where $\odot$ means masking, which indicates only the foreground pixels are applied.
The IUV loss focuses solely on supervising the 2D information, neglecting the depth information essential for capturing distorted details. To address this limitation, we introduce an additional depth map loss. We generate ground truth depth maps by rendering them online. A 2D convolution is used to predict the depth map based on the feature map $\varPhi_{3}$. The loss is computed as follows:
where we use the reciprocal of the depth for normalization.
3 Experiment
3.1 Implementation Details
Pixel-aligned Feature Extractor We follow the feature extraction method of [54] but use different sampling points and backbones. When a cropped image $I\in\mathbb{R}^{H_{c}\times W_{c}}$ is inputted into the encoder, it generates an initial feature map $\varPhi_{0}\in\mathbb{R}^{C\times H_{0}\times W_{0}}$. Subsequently, this feature map $\varPhi_{0}$ undergoes further processing through a series of deconvolution layers to yield high-resolution features $\left\{\varPhi_{i}\in\mathbb{R}^{C_{i}\times H_{i}\times W_{i}}\right\}_{i=1}^{3}$. For the first high-resolution feature map $\varPhi_{1}$, a $H_{g}\times W_{g}$ grid is defined to sample finer point-wise features. But, for the second and third feature maps, we project vertices of the mesh obtained from the previous step onto the feature map to sample point-wise features. In summary, we extract spatial features based on the following formula.
Table 1: This table illustrates the changes in the size of input images and features when using different backbones (ResNet50 or VitPose-B) and employing different downsampling methods (sparse vertices or markers). As for what symbols mean, please refer to Section 3.1.
$H_{c}\times W_{c}$ 224$\times$224 224$\times$224 256$\times$192 256$\times$192
[…]
$C_{i}^{fc}$ 5 32 5 32
$H_{g}\times W_{g}$ 21$\times$21 8$\times$8 24$\times$18 9$\times$7
$\{C_{i}^{s}\}_{i=1}^{3}$2205,2155,2155} {2048,2144,2144} {2160,2155,2155} {2016,2144,2144}
[…]
where $P_{i}$ represents the projected 2D point for sampling, $bs(\cdot)$ denotes bilinear sampling, and $fc(\cdot)$ refers to FC layers for reducing channel size. Finally, $\oplus$ signifies the concatenation of individual point-wise features $\left\{\phi_{i,n}^{p,fc}\in\mathbb{R}^{C_{i}^{fc}}\right\}_{n=1}^{N_{i}}$ into a one-dimensional feature $\phi_{i}^{s}\in\mathbb{R}^{C_{i}^{s}}$, where $C_{i}^{s}=N_{i}\times C_{i}^{fc}$. $fc$ indicates the point-wise features has been through $fc(\cdot)$ for channel reduction and $N_{i}$ is the number of points for downsampling.
Hybrid Regression Module For the spatial features $\{\phi_{i}^{s}\}_{i=1}^{3}$, we have corresponding regression models. We utilize parameter regressors for the first two spatial features to obtain the initial meshes. These regressors take spatial features $\phi_{i}^{s}$, global information $glob\_info$, and former SMPL parameters $\varTheta_{i-1}$ as inputs and output SMPL parameters $\varTheta_{i}$. The parameter regressors consist of several FC layers and iteratively produce the final result like
Note $\varTheta_{0}$ is set as the mean $\overline{\varTheta}$ following [14]. As for the final regressor, we have two versions in total. The first version still involves the parameter regressor. The second version utilizes a vertex regressor to obtain a finer human body mesh. After obtaining a coarse human mesh using the first two regressors, we introduce a model-free vertex regressor based on the transformer. This vertex regressor takes spatial features $\phi_{3}^{s}$ and point-wise features $\left\{\phi_{3,n}^{p}\in\mathbb{R}^{C_{i}}\right\}_{n=1}^{N_{i}}$ as input. It’s important to note that no "$fc$" exists in this case, indicating that these features do not undergo channel reduction. To reduce network complexity and adapt to the SMPL model, instead of directly recovering the 6890 vertices defined by SMPL, we recover 431 sparse vertices and gradually restore them to 6890 vertices using two linear layers.
We train three models in total for evaluation. Two of them utilize a hybrid regression module, using 431 sparse vertices for downsampling. The distinction of these two models lies in using different backbones, including ResNet50 [11] and VitPose-B [48]. Additionally, we train an extra model using a pure parameter regression module, employing 67 mocap markers for downsampling, with the backbone being VitPose-B. For training, we utilize datasets such as Human 3.6m [12], 3DPW [42], MPII3D [31], MPII [1], AGORA [35], COCO [28], SPEC-SYN [19], and HuMMan [3]. Some of the datasets are augmented with pseudo-labels generated by CamCalib, EFT [13], or CLIFF [25]. Please refer to Section 3.2 for detailed information about the datasets used. Our model can simultaneously provide results in camera and world coordinate. But if results in world coordinate are not needed or camera rotation is known, the CamCalib and OrientCorrect can be discarded. The focal length predictor can also be discarded if mesh projection is not required.
We experiment with different backbones and downsampling points to validate our method and explore the optimal model structure. As a result, there are corresponding variations in the sizes of input images and features. We summarize these variations and create Table 1 to provide an overview of these changes. The training process is conducted with a batch size of 128 and a fixed learning rate of 0.00005 on three RTX 3090 GPUs until convergence, taking approximately one week. Following [54, 18], we apply standard data augmentation techniques such as rotation, flipping, scaling, translation, noise, and occlusion. As for hyperparameters, $\lambda_{1}=\lambda_{2}=\lambda_{3}=300$, $\lambda_{4}=\lambda_{5}=\lambda_{6}=15$, $\lambda_{7}=0.06$, $\lambda_{8}=60$, $\lambda_{9}=0.03$ and $\lambda_{10}=\lambda_{11}=\lambda_{12}=0.05$.
3.2 About Datasets
In our experiments, we utilize the following datasets: Human 3.6M [12], 3DPW [42], MPII [1], COCO [28], MPI-INF-3D [31], AVA  [9], AIC [46], Insta [15], AGORA [35], SPEC-SYN [19], SPEC-MTP [19], and HuMMan [3]. We conduct experiments on the distorted datasets AGORA, HuMMan, and SPEC-MTP to evaluate our model. Furthermore, following traditional practices[45, 54], we also evaluate our model on 3DPW, Human 3.6M, and MPII-INF-3D. We employ AGORA’s evaluation platform as a standard for ablation experiments due to its diverse scenes, characters, and shooting methods. We believe that the evaluation results obtained on AGORA are more representative. For human recovery in world space, We additionally perform an ablation study on SPEC-SYN to validate the robustness of our method. Following are details about each dataset.
3DPW’ [42] full name is "3D Human in the Wild." It is an outdoor dataset that captures actor information using motion capture devices. This dataset is known for its difficulty and has been widely used by researchers to demonstrate the robustness of their methods. Following past conventions, we train two models, one without 3DPW for training and another using 3DPW for training, to evaluate our model’s robustness.
Human 3.6M [12] is widely used as a standard dataset to assess the estimation of 3D human pose and the recovery of human mesh. To reduce data redundancy, the original videos are down-sampled from 50 to 10 fps, resulting in 312,188 frames available for training and 26859 frames for evaluation. Following established protocols, we utilize five subjects (S1, S5, S6, S7, S8) for training and two subjects (S9, S11) for evaluation.
Table 2: Comparison with SOTA methods on distorted datasets in camera coordinate. Some of the evaluation results on HuMMan [3] and SPEC-MTP [19] are sourced from  [43], while the evaluation results on AGORA [35] are obtained from its official website https://agora-evaluation.is.tuebingen.mpg.de/. Including the following tables, $*{}^{*}$ represents backbone is ResNet50 [11], ${}^{\dagger}$ means VitPose-B [48, 49], and "P" denotes the pure parameter module. Note we do not perform any fine-tuning on SPEC-MTP.